Hey ive got a problem about system of equations involving quadratic equations:
1.x+y=16
xy=55
2.x-y=3
x^2+y^2=185
3.y^2+xy=-6
x^2+xy=10
4.x^2-y^2=24
xy=35
5.a-b=-1
ab=14
Pls solve it for me.. The main objective is to find the value of x and y on the given.. Tnx --- Jcob

Answer by Alan3354(69443)

1.x+y=16
xy=55
Use substitution.
y = 16-x
Sub for y in the 2nd eqn
=================
2.x-y=3
x^2+y^2=185
Similar to #1
===================
3.y^2+xy=-6
x^2+xy=10
I'll do later.
==============
4.x^2-y^2=24
xy=35
I'll do later.
=====================
5.a-b=-1
ab=14
Similar to #1
=========================================
You work on the 3, email via the TY note for help or to check your work.